Assume that V is constant, M grows at an annual rate of 5%, Y at a rate of

2% and r= 4%.

a)

What will be the nominal interest rate?

b) How will the nominal interest rate change when the CB increases the

growth rate of money supply by 2 percentage points?

c) Assume that the growth rate of Y drops to 1%.

• How will the inflation rate change?

• What must the CB do to ensure that the inflation rate does not change?


1
Expert's answer
2022-03-17T09:43:16-0400

If  V is constant, M is growing 5% per year, Y is growing 2% per year, and r = 4.

a)  Nominal Interest rate= real interest rate + inflation rate

We first find "\pi = 5 -2"

"\pi=3"

Therefore, "i = r + \pi"

"= 4 + 3"

"= 7"

b)

"\Delta i = 2"

It is just the same as the increase in the money growth rate.

"\u03c0=5\u22122"

"\u03c0=3"

Hence, "i = r + \pi"

"=4+3"

"= 7"

c)

The central bank does nothing "\Delta \pi =1". For the central bank to prevent inflation from rising, it must reduce the money growth rate by "1\%" point per year